### v1

@@ Acknowledgements.
AG, MAA, and MO acknowledge support from grant PID2021-127495NB-I00, funded by MCIN/AEI/10.13039/501100011033 and by the European Union “NextGenerationEU” as well as “ESF Investing in your future”. Additionally, they acknowledge support from the Astrophysics and High Energy Physics programme of the Generalitat Valenciana ASFAE/2022/026 funded by MCIN and the European Union NextGenerationEU (PRTR-C17.I1) as well as support from the Prometeo excellence programme grant CIPROM/2022/13 funded by the Generalitat Valenciana. M.O. acknowledgess the support of the Spanish Ramon y Cajal programme (2018-024938-I). GM and SE have received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 833925, project STAREX). RH acknowledges support from the World Premier International Research Centre Initiative (WPI Initiative), MEXT, Japan, the IReNA AccelNet Network of Networks (National Science Foundation, Grant No. OISE-1927130), the European Union’s Horizon 2020 research and innovation programme (ChETEC-INFRA, Grant No. 101008324) and the Wolfson Foundation.
Appendix A The ”Dichte” EoS
@@
The so called Dichte Equation of State — which up until now was the only EoS implemented in GENEC — considers stellar matter as a mixture of perfect ion gas, free electrons and radiation. It further takes into account partial ionisation of the atoms and partial degeneracy of the electrons. In this section we explain in detail the functioning of this EoS and notably the manner in which it takes into account partial ionisation in the stellar envelope. We recall that in the latest models presented in this paper the T-EoS is only used where the condition of Eq. 1 holds, otherwise we resort to the D-EoS.
@@
where $E$ is the number of electrons, per atom, released by ionisation. In the case of partial ionisation, we can find the value of $E$ using the Saha equations. This regime is only reached in the stellar envelopes and is applied only to a selection of the lighter elements common in this region, (i.e H, He, C, O, Ne and Mg), the physics of this module was developed in Schaller et al. (1992). Here, to illustrate the method and for the purpose of simplicity, we explicitly write the equations for the case where only hydrogen and helium are considered. The fraction of the number of atoms of hydrogen and helium present in the mixture are,
@@
With this set of equations one may solve iteratively the Saha equations (35)-(37) and determine the value of $E$, and thus the electron pressure for partial ionisation. Note that we consider complete ionisation for hydrogen and helium, ($x_{1}=1$, $x_{2}=0$ and $x_{3}=1$ ) when the ionisation pressure is sufficiently large so that the Saha equations are no longer valid. In practice, we assume complete ionisation if (Kippenhahn et al. 2013, § 14.6)
@@
where $\beta=P_{\rm g}/P$. In this case the electron pressure is computed by taking all species as totally ionised not just hydrogen and helium and summing their contributions to compute $E$ and $P_{\rm e}$. This is accurate until we reach the regime of partial degeneracy where then the electron pressure is treated differently, i.e., when,
@@
which roughly corresponds to an electron degeneracy parameter $\psi\lesssim-4$ (Hofmeister et al. 1964). Under the conditions of full ionisation, the electron pressure is computed using suitable approximations of the Fermi-Dirac integrals, including the both the non-relativistic and relativistic regimes. Given an input pressure and temperature, $(P,T)$, the D-EoS assumes that,
@@
As in the regime of partial ionization, the gas pressure results from the contribution of the ions and of the free electrons,

### v2

@@ Acknowledgements.
AG, MR, MAA, and MO acknowledge support from grant PID2021-127495NB-I00, funded by MCIN/AEI/10.13039/501100011033 and by the European Union “NextGenerationEU” as well as “ESF Investing in your future”. Additionally, they acknowledge support from the Astrophysics and High Energy Physics programme of the Generalitat Valenciana ASFAE/2022/026 funded by MCIN and the European Union NextGenerationEU (PRTR-C17.I1) as well as support from the Prometeo excellence programme grant CIPROM/2022/13 funded by the Generalitat Valenciana. M.O. acknowledges the support of the Spanish Ramon y Cajal programme (2018-024938-I). MR acknowledges support from the Spanish Juan de la Cierva programme (FJC2021-046688-I). GM and SE have received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 833925, project STAREX). RH acknowledges support from the World Premier International Research Centre Initiative (WPI Initiative), MEXT, Japan, the IReNA AccelNet Network of Networks (National Science Foundation, Grant No. OISE-1927130), the European Union’s Horizon 2020 research and innovation programme (ChETEC-INFRA, Grant No. 101008324) and the Wolfson Foundation.
Appendix A The GENEC code
The stellar evolution code GENEC solves the internal structure equations as described in the appendix A of Meynet & Maeder (1997) and where the numerical methods employed are detailed in Kippenhahn et al. (1967). A complete description of the code (without the improvements included in this paper) can be found in Eggenberger et al. (2008). The discretisation of the equations follows the prescription of Sugimoto (1970), but does not include any acceleration term (i.e. they assume hydrostatic equilibrium). The nuclear reaction network during the main sequence is essentially the same as employed in Meynet & Maeder (2003), and up to carbon burning it follows closely the implementation done in the PhD thesis Hirschi (2004), with the updates in reaction rates included in Ekström et al. (2012). Modifications in the late phases of evolution are included here (Sect. 4). The treatment of convection and diffusion can be found, for example in Ekström et al. (2012) and references therein. Rotation is implemented following the basic prescription of Meynet & Maeder (1997), where the structure equations are averaged out on isobars and a Roche-approximation (see, e.g. Maeder 2009,  § 2.1) is employed. GENEC includes the action of meridional circulation currents, employing an advective treatment (instead of diffusive; e.g. Maeder & Zahn 1998).
Next, we present briefly some further details about the numerical efficiency of the GENEC models computed in this paper. In GENEC the typical number of mass shells during evolution is $\approx$ 1500, which correspond to an average mass resolution of $\delta m=5\times 10^{-3}M_{\odot}$. This is to be compared with the MESA models which use an average of 14000 mass shells corresponding to an average mass resolution of $\delta m=5\times 10^{-4}M_{\odot}$. We note that the number of mass shells in GENEC adapts dynamically during the evolution, adding shells where higher resolution is needed to capture strong gradients, or removing them where such resolution is no longer needed. Layers may be added or removed depending on whether various criteria are met. For example, if the variation in pressure, radius or luminosity between two cells is very small then these cells are fused, but, on the contrary, if the relative variations of any of these quantities is too large, cells are added.
The typical number of time-steps used by a MESA model is on the order of 20000, whereas for a GENEC model this can be up to 6-7 times more. Noteworthy in GENEC simulations, more than 60% of the steps correspond to the evolution posterior to oxygen core exhaustion, highlighting the numerical difficulty of this late phase of evolution. The small time steps needed for the evolution in that phase are produced by the stiffness in the nuclear reaction network. We point out that in the solution strategy of GENEC, the temperature is assumed constant during a burning sub-step. After it, temperature (and the rest of the variables in the Henyey iteration) are updated. Further iterations of alternating sub-steps of burning and of the Henyey solver are performed until convergence is found and a time step is completed. During the late phases of evolution, this strategy limits significantly the time step size for stability, as well as for accuracy. We remark that the small steps taken in GENEC are also enforced to maintain the sum of mass fractions as close to one as possible, with an absolute tolerance $1-\sum_{i}X_{i}<10^{-5}$. That is enforced without renormalising mass fractions after any burning step. In contrast, MESA, solves the nuclear network along with the temperature, allowing for renormalisation of the isotopic abundances (potentially numerically modifying these abundances during the normalisation), providing longer time steps over which the solution is stable. Future versions of GENEC may consider this solution strategy to improve the numerical performance of the code.
Appendix B The Dichte EoS
@@
The Dichte Equation of State, which up until now was the only EoS implemented in GENEC, considers stellar matter as a mixture of perfect ion gas, free electrons and radiation. It further takes into account partial ionisation of the atoms and partial degeneracy of the electrons. In this section we explain in detail the functioning of this EoS and notably the manner in which it takes into account partial ionisation in the stellar envelope. We recall that in the latest models presented in this paper the T-EoS is only used where the condition of Eq. 1 holds, otherwise we resort to the D-EoS.
@@
where $E$ is the number of electrons, per atom, released by ionisation. In the case of partial ionisation, we can find the value of $E$ using the Saha equations. This regime is only reached in the stellar envelopes and is applied only to a selection of the lighter elements common in this region, (i.e H, He, C, O, Ne, and Mg), the physics of this module was developed in Schaller et al. (1992). Here, to illustrate the method and for the purpose of simplicity, we explicitly write the equations for the case where only hydrogen and helium are considered. The fraction of the number of atoms of hydrogen and helium present in the mixture are,
@@
With this set of equations one may solve iteratively the Saha equations (36)-(38) and determine the value of $E$, and thus the electron pressure for partial ionisation. We note that we consider complete ionisation for hydrogen and helium, ($x_{1}=1$, $x_{2}=0$ and $x_{3}=1$ ) when the ionisation pressure is sufficiently large so that the Saha equations are no longer valid. In practice, we assume complete ionisation if (Kippenhahn et al. 2013, § 14.6)
@@
where $\beta=P_{\rm g}/P$. In this case the electron pressure is computed by taking all species as totally ionised not just hydrogen and helium and summing their contributions to compute $E$ and $P_{\rm e}$. This is accurate until we reach the regime of partial degeneracy where then the electron pressure is treated differently, that is when
@@
which roughly corresponds to an electron degeneracy parameter $\psi=\lesssim-4$ (Hofmeister et al. 1964). Under the conditions of full ionisation, the electron pressure is computed using suitable approximations of the Fermi-Dirac integrals, including the both the non-relativistic and relativistic regimes. Given an input pressure and temperature, $(P,T)$, the D-EoS assumes that,
@@
As in the regime of partial ionisation, the gas pressure results from the contribution of the ions and of the free electrons,
@@
Appendix C Matching between the D-EoS and the T-EoS
The transition between the D-EoS and the T-EoS, is done when condition (1) is satisfied.
Figure 16: Evolution of the central sound speed, $c_{s}=\sqrt{\partial_{\rho}P}$, as a function of central density. The matching point between the two EoSs is at the central density value $\log(\rho_{c})=2.8$, except for the lower threshold model which changes EoS at $\log(\rho_{c})=1.8$.
To demonstrate that the matching point does not have spurious effects on the models, we plot the central value of the sound speed and of $\delta$ (Eq. 51) as a function of central density. In Figs. 16 and 17, we can see the smooth transition of these thermodynamic derivatives through the density threshold located at $\log(\rho_{\rm matching})=2.8$. We note that for the $25M_{\odot}$ model (black line) there is a sharpness close to the threshold, this is not due to the EoS change but rather the evolutionary stage it has reached at that point. To demonstrate this, we show how the evolution of the same $25M_{\odot}$ model but using a lower density threshold ($\log(\rho_{\rm matching})=1.8$; dashed yellow line) overlaps the evolution of the former model with the standard transition threshold in the case of the sound speed (Fig. 16). The evolution of $\delta$ (Fig. 17) shows a tiny difference after crossing the lower density threshold, which tends to disappear once the central density rises by one order of magnitude. Thus, the matching of the two equations of state here is robust and does not produce numerical artefacts that could be troublesome for the computation.
Figure 17: Evolution of the thermodynamic derivative $\delta$, defined in Eq. 51, computed at the centre of the star as a function of central density. The matching point between the two EoSs is at the central density value $\log(\rho_{c})=2.8$, except for the lower threshold model, which changes EoS at $\log(\rho_{c})=1.8$.